OpenStudy (anonymous):

The line through (4,2) parallel to y=-5/7x +1. Show me how to solve using slope formula? Thank you. No answer needed, i have them, just studying. :)

OpenStudy (anonymous):

i will just take it as it says slope

OpenStudy (anonymous):

if you want a line parallel just give it the same slope of that as the same to y=-5/7x +1

OpenStudy (anonymous):

slope formula y1 - y2 m = --------- x1 - x2 your problem " line through (4,2) " line through (4,2) implies that it will cross this point so this will be one of the points "parallel" implies it will never will always have the same distance between line and will never intersect therefore it must have same slope as original problem (4)1 - y2 5/7 = -------- ( 2 )1 - x2 parenthesis here do not mean to multiply they mean substitute for this case

OpenStudy (anonymous):

u good?

OpenStudy (anonymous):

i'm not sure, i thought there was only one formula

OpenStudy (anonymous):

i am only using one formula right now

OpenStudy (anonymous):

it is slope formula the other one you see there is your question in the problem that is posted

OpenStudy (anonymous):

isnt that midpoint formula?

OpenStudy (anonymous):

no it is not

OpenStudy (anonymous):

mid point you have divide 2

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but do you understand the logic to this work i did?

OpenStudy (anonymous):

hypothetical example  parallel lines must have same slope same distance from each other to satisfy as parallel lines

OpenStudy (anonymous):

thank you